The length of a rectangle is 4 more than its width. The area is 60 sq. yards. Find the dimensions.
Answer by mananth(16946)

let width be w
length = w+4
..
area =60
w(w+4)=60
w^2+4w=60
w^2+4w-60=0
(w+10)(w-6)=0
w= 6 width
length = 10 yards
...
m.ananth@hotmail.ca